How can you tell if a parabola opens left/right or up/down such as for this question.
The foucs (-3,5) diretrix at x= 1 --> write it in standard form and tell whether it opens right/left or up/dpwn.

Answer by stanbon(48569)

Draw the picture of the focus and the directrix on
a coordinate system. Notice the focus is "4" to the
left of the directrix. So the parabola opens to the
left. The vertex is half way between the focus and
the directrix so the vertex is at (-1,5)
The standard form for your parabola is as follows:
(y-k)^2 = 4p(x-h)
h=-1, k=5,
p is the distance from the vertex to the focus = -2
The equation is (y-5)^2 = -8(x+1)
I think I am defining "p" correctly but you might check
it out in your textbook.
Cheers,
stan H.